Curriculum Vitae for Oliver Goodman

Personal Details

Dr. Oliver A. Goodman  The Geometry Center
Phone: (612) 331-7651  University of Minnesota
Email: oag@geom.umn.edu  1300 South 2nd St., Suite 500
WWW: http://www.geom.umn.edu/∼oag/  Minneapolis, MN 55454

Nationality: British Education

1989 PhD. in mathematics from Warwick University, Coventry, England, supervised by Professor Epstein. Thesis title: “Metrized Laminations and Quasisymmetric Maps.” 1987 MSc. in mathematics (with distinction) from Warwick University, Coventry, England. 1986 BA. in mathematics (1st class) from St Catherine’s College, Oxford, England.

Professional Experience Jan 1995-present: European Union fellowship at the University of Bielefeld. Working with D.Huson, author of the RepTiles computer program, on hyperbolic and automatic groups. Studying potentially computable problems in the homology of groups. Oct-Dec 1994: Visiting position at the University of Melbourne. Worked with Hodgson and Neumann implementing exact arithmetic in the 3-manifold program SnapPea for a study of arithmetic Kleinian groups. Mar 1992-Oct 1994: Postdoc at the Geometry Center, University of Minnesota. Worked mainly on mathematical software; in particular, computation and geometric visualization in hyperbolic geometry. Studied related theoretical issues from a mathematical point of view. Taught one quarter of pre-calculus. Jan-Mar 1992: Worked for the Geometry Center at Warwick University writing software to compute circle packings. Jan 1990-Dec 1991: Programmer/Analyst at IBM UK. Gained much useful programming experience.

Research and software developed “On the tight span of a 2-dimensional spherical graph”: A study of the tight span construction which provides a route to the computation of the homology a group. In preparation. “A constructive proof of Andre’ev’s theorem in the compact case”: Proof of Andre’ev’s theorem and details of an algorithm for carrying out the construction. In preparation. Hyperbolic.m: Mathematica package for carrying out computations in n-dimensional hyperbolic geometry and displaying 2- and 3-dimensional results visually in a variety of models. Available by ftp from geom.umn.edu

as pub/software/hyperbolic.tar.Z.

CirclePack.m: Mathematica package to compute circle packings on geometric surfaces. Available by ftp from

geom.umn.edu as pub/software/circlepack.tar.Z.

Geomview conformal model: Designed and carried out the initial implementation of Geomview’s Poincaré ball model view of hyperbolic 3-space. Geomview is available from geom.umn.edu as pub/software/geomview. Mathpad: An Iris based interactive graphics tool for Mathematica, designed to overcome the noninteractive nature of Mathematica graphics. To be released.

Conferences and Visits Oct 1995: Talked about my work with Hodgson and Neumann at AMS meeting in Boston. Sep 1993: Talked on Andre’ev’s theorem and circle packings at North Illinois University. Oct 1992: Presented Hyperbolic.m at a conference on geometric visualization at the MSRI in Berkeley.